if x =arctan(-1) = -pi/4
what is x = arctan (4) = ??? radians
what is arctan (4) in radians?

Answer by stanbon(75887)

if x =arctan(-1) = -pi/4
Note: x must be in the 4th Quadrant
------------------------------
what is x = arctan (4) = ??? radians
x is the angle whose tangent is 4:
tan^-1(4) = 75.96 degrees or 255.96 degrees
-------
what is arctan (4) in radians?
x = (75.96)(pi/180) = 0.4220pi
or
x = (255.96)(pi/180) = 1.4220pi
===============
Cheers,
Stan H.